According to a Pew Research Center study, in May 2011, 30% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 280 community college students at random and finds that 90 of them have a smart phone. Then in testing the hypotheses:

H0: p = 0.3 versus

Ha: p > 0.3,

what is the test statistic?

z = . (Please round your answer to two decimal places.)
Answer #1

Sample proportion = 90 / 280 = 0.3214

Test statistics

z = - p / sqrt( p (1 - p) / n)

= 0.3214 - 0.30 / sqrt( 0.30 * 0.70 / 280)

= 0.78